# Bayesian Optimization for Sample-Efficient Policy Improvement in Robotic Manipulation

### v1

@@ II Related Work
Most closely related to the approach that we propose are [13, 14, 15]. Englert et al. [13] propose using BOpt to identify a low-dimensional task mapping for a constrained optimization problem which they initialize from a single demonstration. They continue to sample possible mappings and finally use their motion problem with the uncovered mapping to perform control to solve their tasks. With their approach, they can optimize policies for manipulating articulated objects. They reward their agent with the negative forces exerted during the interactions. While their data efficiency is impressive, the space of their mapping is extremely low-dimensional with no more than three dimensions. Johannsmeier et al. [14] model a manipulation skill as a chain of primitive controllers with learnable parameters for desired contact force and superimposed oscillations. They study the suitability of different optimizers for three example tasks and find BOpt to be sufficient, though it is outperformed by an evolutionary strategy. The low success of BOpt might be explainable with their 30-dimensional parameter space and rather large parameter range. In their setups, they derive a cost from the execution duration of a robotic skill. Wu et al. [15] follow up the efforts of [14, 31] by again employing BOpt to learn parameters for motion primitives. Their key change is encoding the demonstration trajectories as a GMM and incorporating the probability of the current trajectory under this distribution in the optimizer’s objective function, enforcing a similarity of generated trajectories to the original demonstrations. This similarity function is the main reward signal for the optimizer, with only minor influence given to task success.
@@ IV-A GMM
In this work, we examine the challenge of improving an initial robotic motion policy $\pi_{\theta}$ trained from a set of demonstrated trajectories $\Xi_{D}$. Our approach assumes motion to be driven by a dynamical system of the form $f_{\theta}(s)=\dot{s}$, encoded in a GMMs as parameterization of $\pi_{\theta}$. Here $\dot{s}$ denotes the first-order derivative of the observable state $\mathbf{s}$. We follow the Dynamical System definition of [8, 27]. Given an observable system state $\mathbf{s}\in\mathbb{R}^{m}$ a GMM models the dynamics of this system as $K$ components, weighted by $\mathbf{\omega}\in[0,1]^{K}$ with $\left\lVert\mathbf{\omega}\right\rVert=1$, each of which consists of a $2m$-dimensional mean $\mu^{k}$, and a corresponding covariance matrix $\Sigma^{k}$, with
@@ IV-B GMM Parameterization
$\Sigma_{i}^{k}=\mathbf{Q}\mathbf{\Lambda}\operatorname{\text{diag}}(\mathbf{\Delta\theta}_{\Sigma^{k},i})\mathbf{Q}^{-1},$ (6)
@@ IV-C Bayesian Optimization
Bayesian Optimization (BOpt) is the state-of-the-art technique for hyperparameter tuning in automated machine learning tasks. In this domain, its problem is formulated as finding a vector of hyperparameters $\mathbf{x}\in A$ which maximizes model performance as $\max_{x\in A}f(x)$, where $A$ is typically a bounded hypercube in $\mathbb{R}^{n}$. Unlike model-free techniques such as grid search and random search, BOpt algorithms build a probabilistic surrogate function $p\left(y\;\middle|\;\mathbf{x},D\right)$, where $D$ is a dataset of evaluated hyperparameter samples, as defined in Sec. III and $\mathbf{x}$ is a newly generated hyperparameter sample. The function’s probabilistic nature allows the optimizer to explore the parameter space according to the expected value as estimated by the surrogate [32, 33]. A common surrogate implementation is to use Gaussian Processes (GP), however, these lend themselves mostly to lower-dimensional parameter spaces and are costly to evaluate on large datasets [32]. Random forests (RF) are a common alternative to using GPs in BOpt [32, 33]. By using multiple random trees as regressors and averaging their output, RFs are able to provide both an expectation and uncertainty estimate of a given $\mathbf{x}$, while being much more time and space-efficient than GPs. As our parameter space is quite large and our data becomes plentiful over time, we use RFs in this work.
@@ V-B Experiments in Simulation
We collect 10 demonstrations in each scenario and fit a GMM for each setting. As the number $k$ of Gaussians used has an impact on the performance of the model while also scaling the number of parameters to optimize, we perform evaluations with $k\in\left\{3,5,7\right\}$. We give both SAC-GMM and the Bayesian optimizer update ranges of $\Delta\theta_{\omega}\in[-0.1,0.1]$, $\Delta\theta_{\mu}\in[-0.05,0.05]$, and $\Delta\theta_{\Sigma}\in[-0.1,0.1]$. We use the Bayesian Optimization implementation from SMAC3 [35], a collection of mature gradient-free optimizers for black-box optimization. Specifically, we use the hyperparameter optimizer with logarithmic expected improvement as an acquisition function. In addition, we set SAC-GMM’s learning rate as $2\cdot 10^{-3}$. The baseline performances of these models are presented in Fig. 4.
@@
Mean
SAC $\mathbf{88.4\%}$ $258.3$ $\mathbf{95.3\%}$ $125.0$ $\mathbf{94.2\%}$ $183.3$ $\mathbf{88.1\%}$ $258.3$ $\mathbf{84.1\%}$ $283.3$ $\mathbf{90.0\%}$ $221.6$

### v2

@@ II Related Work
Using Bayesian Optimization methods for policy improvement is not a new approach. BOpt has been used successfully in finding threshold parameters for walking gaits [31, 32] as well as efficient impact-compensating arm movements for balancing robots [33]. In [34] the authors demonstrate sample-efficient learning of a robotic policy in simple scenarios by learning function of the task dynamics as a Gaussian Process (GP). A similar technique is employed by [35], who additionally use a linear projection to reduce the number of policy parameters to. More flexible is the approach presented in [36] which uses an auto-encoding scheme to form a low-dimensional parameter space, trained on simulated trajectories. BOpt has also been used to decide sampling a simulation or a real world execution [37], reducing the number of expensive real world examples needed to improve a policy. Different to our proposed method all of these methods require a dense reward signal. Using binary signals in combination with BOpt has been studied as well, though less throughly. In [38], the authors formulate an approach to maximizing a stochastic binary reward by using a GP to model the parameters of a binomial distribution. While they do demonstrate that this model can be used to exploit such a function efficiently, we use a different approach to integrating binary feedback, as this model only lends itself to univariate scenarios.
Most closely related to the approach that we propose are [13, 14, 15]. Englert et al. [13] propose using BOpt to identify a low-dimensional task mapping for a constrained optimization problem which they initialize from a single demonstration. They continue to sample possible mappings and finally use their motion problem with the uncovered mapping to perform control to solve their tasks. With their approach, they can optimize policies for manipulating articulated objects. They reward their agent with the negative forces exerted during the interactions. While their data efficiency is impressive, the space of their mapping is extremely low-dimensional with no more than three dimensions. Johannsmeier et al. [14] model a manipulation skill as a chain of primitive controllers with learnable parameters for desired contact force and superimposed oscillations. They study the suitability of different optimizers for three example tasks and find BOpt to be sufficient, though it is outperformed by an evolutionary strategy. The low success of BOpt might be explainable with their 30-dimensional parameter space and rather large parameter range. In their setups, they derive a cost from the execution duration of a robotic skill. Wu et al. [15] follow up the efforts of [14, 39] by again employing BOpt to learn parameters for motion primitives. Their key change is encoding the demonstration trajectories as a GMM and incorporating the probability of the current trajectory under this distribution in the optimizer’s objective function, enforcing a similarity of generated trajectories to the original demonstrations. This similarity function is the main reward signal for the optimizer, with only minor influence given to task success. Otto, Celik,  et al. [40, 41] phrase a policy improvement problem on DMPs similar to ours, but use a deep function approximator which leads to a significant data requirement.
@@ IV-A GMM
In this work, we examine the challenge of improving an initial robotic motion policy $\pi_{\theta}$ trained from a set of demonstrated trajectories $\Xi_{D}$. Our approach assumes motion to be driven by a dynamical system of the form $f_{\theta}(s)=\dot{s}$, encoded in a GMMs as parameterization of $\pi_{\theta}$. Here $\dot{s}$ denotes the first-order derivative of the observable state $\mathbf{s}$. We follow the Dynamical System definition of [8, 27]. Given an observable system state $\mathbf{s}\in\mathbb{R}^{m}$, in our case the position of the robot’s endeffector relative to a frame of reference, a GMM models the dynamics of this system as $K$ components, weighted by $\mathbf{\omega}\in[0,1]^{K}$ with $\left\lVert\mathbf{\omega}\right\rVert=1$, each of which consists of a $2m$-dimensional mean $\mu^{k}$, and a corresponding covariance matrix $\Sigma^{k}$, with
@@ IV-B GMM Parameterization
$\Sigma_{x,i}^{k}=\mathbf{Q}\mathbf{\Lambda}\operatorname{\text{diag}}(\mathbf{\Delta\theta}_{\Sigma^{k},i})\mathbf{Q}^{-1},$ (6)
@@
We also consider a rank-1 vector-base update for $\Sigma_{i}^{k}$ based on a singular value decomposition and rank reduction. We find this update method to be less stable than the ones presented above, but we include a brief description of it for the interested reader in Sec. APPENDIX - Low-rank Covariance Updates.
@@ IV-C Bayesian Optimization
Bayesian Optimization (BOpt) is the state-of-the-art technique for hyperparameter tuning in automated machine learning tasks. In this domain, its problem is formulated as finding a vector of hyperparameters $\mathbf{x}\in A$ which maximizes model performance as $\max_{x\in A}f(x)$, where $A$ is typically a bounded hypercube in $\mathbb{R}^{n}$. Unlike model-free techniques such as grid search and random search, BOpt algorithms build a probabilistic surrogate function $p\left(y\;\middle|\;\mathbf{x},D\right)$, where $D$ is a dataset of evaluated hyperparameter samples, as defined in Sec. III and $\mathbf{x}$ is a newly generated hyperparameter sample. The function’s probabilistic nature allows the optimizer to explore the parameter space according to the expected value as estimated by the surrogate [42, 43]. A common surrogate implementation is to use Gaussian Processes (GP), however, these lend themselves mostly to lower-dimensional parameter spaces and are costly to evaluate on large datasets [42]. Random forests (RF) are a common alternative to using GPs in BOpt [42, 43]. By using multiple random trees as regressors and averaging their output, RFs are able to provide both an expectation and uncertainty estimate of a given $\mathbf{x}$, while being much more time and space-efficient than GPs. As our parameter space is quite large with up to $70$ parameters, and our data becomes plentiful over time, we use RFs in this work.
@@ V-B Experiments in Simulation
We collect 10 demonstrations in each scenario and fit a GMM for each setting. As the number $k$ of Gaussians used has an impact on the performance of the model while also scaling the number of parameters to optimize, we perform evaluations with $k\in\left\{3,5,7\right\}$, which yield at most $30,50$ and $70$ parameters respectively. Throughout our experiments we use $j=8$ evaluation episodes. We give both SAC-GMM and the Bayesian optimizer update ranges of
$\Delta\theta_{\omega}$ $\in[-0.1,0.1]$ (8)
$\Delta\theta_{\mu}$ $\in[-0.05,0.05]$
$\Delta\theta_{\Sigma}$ $\in[-0.1,0.1].$
We use the Bayesian Optimization implementation from SMAC3 [45], a collection of mature gradient-free optimizers for black-box optimization. Specifically, we use the hyperparameter optimizer with logarithmic expected improvement as an acquisition function. In addition, we set SAC-GMM’s learning rate as $2\cdot 10^{-3}$. The baseline performances of these models are presented in Fig. 4.
@@
Mean SAC $\mathbf{88.4\%}$ $258.3$ $\mathbf{95.3\%}$ $125.0$ $\mathbf{94.2\%}$ $183.3$ $\mathbf{88.1\%}$ $258.3$ $\mathbf{84.1\%}$ $283.3$ $\mathbf{90.0\%}$ $221.6$
@@ VI Conclusion
APPENDIX - Low-rank Covariance Updates
As additional GMM size $|S|$ update, we examine a rank-1 update scheme. The scheme is based on constructing a low-rank approximation of $\Sigma^{k}$, performing an update under this approximation and reintegrating the update into the full-rank matrix. We do so by extracting $\mathbf{U},\mathbf{s},\mathbf{V}$ from $\Sigma^{k}$ as
$\Sigma^{k}-\operatorname{\text{diag}}(\Sigma^{k})=\mathbf{U}\operatorname{\text{diag}}(\mathbf{s})\mathbf{V}^{T},$ (9)
using singular value decomposition. We use this decomposition to formulate a rank-1 update by only considering the first column $\mathbf{U}_{1},\mathbf{V}_{1}$ of $\mathbf{U},\mathbf{V}$ and the first component $\mathbf{s}_{1}$ of $\mathbf{s}$. Given the $i$-th update vector $\Delta\theta_{\Sigma^{k},i}\in[1-\sigma,1+\sigma]$, we compute the updated matrix $\Sigma_{i}^{k}$ as
$\Sigma_{i}^{k}=\operatorname{\text{diag}}(\Sigma^{k})+\mathbf{U}_{1}\operatorname{\text{diag}}(\mathbf{s}_{1})(\mathbf{V}_{1}\cdot\Delta\theta_{\Sigma^{k},i})^{T}.$ (10)
We investigate this update type in our simulated scenarios with both BOpt-GMM and SAC-GMM. We find that this update type is much less reliable at achieving strong performance with either method and varying number of GMM components, see Tab. III. Upon inspection, we find that the rank-1 update performs well ($\geq 90\%$) in some settings, but fails in others ($\leq 30\%$). We suspect this is due to the required decomposition being executed on $\Sigma^{k}-\operatorname{\text{diag}}(\Sigma^{k})$, which can become numerically unstable as $\Sigma^{k}$ becomes more and more diagonal. However, we need this decomposition to maintain the invertibility of $\Sigma_{i}^{k}$ as required by the GMR inference, see Eq. (4) for details. In addition, this method allows us to maintain the positive determinant of $\Sigma^{k}$, which is required by the function $h^{k}(s_{t})$ also required during inference.
TABLE III: We display the average final success rates achieved by our ablated rank-1 update method. The reported success rates are averages over all simulated scenarios and all numbers of GMM components $k$. While we see in Fig. 4 that BOpt-GMM achieves an average success rate of over $80\%$ independent of the update type and $k$, this is not the case for the low-rank update. Similarly SAC-GMM underperforms with this update. We show the aggregated performance of the other update types in the right-most column.
[…]
Rank-1 Updates
SAC $73.9\%$ $72.0\%$ $73\%$ $88.7\%$
BOpt $74.1\%$ $72.1\%$ $73.1\%$ $87.7\%$